# Generative Diffusion Model (GDM) for Optimization of Wi-Fi Networks

### v1

@@ V Conclusion
In this paper, we introduce a MAC layer access mechanism based on the D3PG algorithm, which enhances network throughput by jointly adjusting the contention window and frame aggregation for each STA. This mechanism addresses the sharp decline in throughput performance caused by increased competition within BSS in dense Wi-Fi scenarios under the existing 802.11 Wi-Fi standard. The throughput improvements with D3PG optimization are significant, showing a 76.4% increase over the baseline 1 and surpassing two other reinforcement learning algorithms baseline 2 and baseline 3 by 13.5% and 10.5%, respectively. Besides, supported by generative diffusion models, which excel at modeling complex data distributions, the D3PG algorithm offers more stable training, faster convergence, and greater flexibility and adaptability. We initially applied GAI to optimize the MAC layer performance of Wi-Fi networks. The results indicate that this approach effectively enhances performance, but the scenarios are relatively simple, and the effectiveness of further applications requires more in-depth study.

### v3

@@ V Conclusion
In this paper, we introduce a MAC layer access mechanism based on the D3PG algorithm, which enhances network throughput by jointly adjusting the contention window and frame aggregation for each STA. This mechanism addresses the sharp decline in throughput performance caused by increased competition within BSS in dense Wi-Fi scenarios. The throughput improvements with D3PG optimization are significant, showing a 76.4% increase over the baseline 1 and surpassing baseline 2 and baseline 3 by 13.5% and 10.5%, respectively. Besides, supported by generative diffusion models, which excel at modeling complex data distributions, the D3PG algorithm offers more stable training, faster convergence, and greater flexibility and adaptability. We initially applied GAI to optimize the MAC layer performance of Wi-Fi networks, and the results indicate that this approach effectively enhances performance.